OpenStudy (hhopke):

What are the domain and x-intercept of: g(x)=25√-24x+27 ??? WILL MEDAL AND FAN!!!

OpenStudy (campbell_st):

so is the question \[g(x) = 25\sqrt{-24x + 27}\]

OpenStudy (hhopke):

yeah- what's the domain (in interval notation), and what is the horizontal intercept (as a coordinate pair)?

OpenStudy (freckles):

horizontal intercept?

OpenStudy (hhopke):

yeah.

OpenStudy (freckles):

I never heard it call that

OpenStudy (hhopke):

or x-intercept?

OpenStudy (freckles):

x-intercept can be found by replacing y with 0 and solving for x

OpenStudy (freckles):

domain just remember numbers under square root thingy must be positive or 0

OpenStudy (campbell_st):

ok... so the x-intercept will also also be the upper limit on the domain you need to look at the part inside the square root... remember, you can't take the square root of a negative so you need to solve \[-24x + 27 \ge 0\] to get the domain the x-intercept, you need to solve \[-24x + 27 = 0\]

OpenStudy (campbell_st):

hope it all makes sense

OpenStudy (hhopke):

'kay, so -24x=-27 -27/-24=x x=1.125 so the x intercept would be (1.125,0) and the domain would be (-oo,1.125)? the domain one sounds weird- I feel like there's some test or something to know whether it's (-oo,1.125) or (1.125,oo) *oo is how I was taught to type infinity

OpenStudy (campbell_st):

well remember, in the domain, you solved \[-24x + 27 \ge 0\] to the domain contains 9/8 or 1.1.125 so just change the bracket and everything else is correct

OpenStudy (hhopke):

Thanks!